OpenStudy (friendzone):

From a standard deck of cards, what is the probability that you draw a diamond, replace it and then draw a club? A) 1 16 B) 1 4 C) 1 2 D) 1

OpenStudy (anonymous):

A: 1/16

OpenStudy (anonymous):

On the first draw, you have a 1/4 chance of drawing a card from 1 of the 4 suits. In other words, you have a 13/52 chance in drawing one of the 13 cards in a suit. 13/52=1/4. If you "replace" the card you just drew (i.e., put it back in the deck), then you'll have a 13/52 chance of drawing another card of a particular suit. Because the two draws are independent, you can multiply the two probabilities, so the probability of getting a diamond and then a club is 13/52*13/52 = 1/4 * 1/4 = 1/16.

OpenStudy (anonymous):

oh you posted this loads of time ha sorry